OpenStudy (jake124):

-12 = f/0.2 A. -2.4 B. 0.24 my answer C. -24 D. 2.4

OpenStudy (mathsucks321):

Its A

OpenStudy (jake124):

thank you !!!!

OpenStudy (mathsucks321):

How did you get 0.24

OpenStudy (3mar):

If you want to solve for f: \[-12=\frac{ f }{ 0.2 }\] \[-12*0.2=\frac{ f }{ 0.2 }*0.2\] \[-2.4=f\]